OpenStudy (anonymous):

The distance from the sun to the nearest star is about 4x10 to the 16th meters. The Milky Way galaxy is roughly a disk of diameter ~ 10 to the 21power meter and thickness ~ 10 to the 19power meter. Find the order of magnitude of a number of stars in the milky way. assume the distance between the sun and our nearest neighbor is typical.

OpenStudy (anonymous):

if typical distance between stars is s, stars per unit volume = 1/(4/3*pi* s ^3 ) Volume of the milky way = pi * r^2 * d , where r is radius and d is thickness of the galaxy. hence estimate of stars in the galaxy = stellar density * volumne = 4.6875 * 10 ^ 11